OpenStudy (anonymous):

Someone please help: Integral help. Integrate: sinx/x^2 dx

OpenStudy (dumbcow):

integration by parts

OpenStudy (anonymous):

Integrate by parts twice,\[\int\limits \frac{dv}{dx}udx=uv-\int\limits v\frac{du}{dx}dx\]Treat sinx as v both times, as the sign will reverse and you can cancel.

OpenStudy (anonymous):

is the answer -sinx/x ?